Solve using substitution method
6m + n = 29
m - 6n = 48

Answer by stanbon(75887)

Solve using substitution method
6m + n = 29
m - 6n = 48
----------------
m = 6n+48
---
Substitute for "m" and solve for "n":
6(6n+48) + n = 29
37n + 288= 29
n = -7
=================
Since m = 6n+48,
m = 6*(-7)+48
m = 6
==================
Cheers,
Stan H.